# Upper Mid-Band Channel Measurements and Characterization at 6.75 GHz FR1(C) and 16.95 GHz FR3 in an Indoor Factory Scenario

### v2

@@ III-C RMS Angular Spread
The RMS AS shows the spatial dispersion of arrival or departure angles in azimuth and elevation planes. RMS AS is evaluated for a spatial lobe threshold of 10 dB below the strongest pointing direction, using procedures detailed in [22]. Table V summarizes the AS observed for both arrival angles at the RX (ASA) and departure angles at the TX (ASD). In LOS, the ASA is observed as 23.71∘ at 6.75 GHz and 14.32∘ at 16.95 GHz. Wide ASAs of 66.38∘ and 50.40∘ are observed in NLOS at 6.75 GHz and 16.95 GHz, respectively. 3GPP models a frequency-dependent LOS ASA with overestimated spread, along with a constant NLOS ASA, as outlined in Table 7.5-6 Part-3: Channel Model Parameters for InF and Table 7.5-11: ZSD and ZOD Offset Parameters for InF in [13]. ASAs are observed to decrease with frequency in both LOS and NLOS. Further, the ASDs are observed as 35.89∘ in LOS and 45.93∘ in NLOS at 6.75 GHz, while ASDs of 22.91∘ in LOS and 34.77∘ in NLOS are observed at 16.95 GHz. As shown in Fig. 4, the CDFs of empirical data align with these trends, while 3GPP predictions suggest a constant ASD of 39∘ in LOS and 38.9∘ in NLOS across frequencies.
@@ IV Conclusion
We presented an in-depth analysis of radio propagation behavior for InF at 6.75 GHz and 16.95 GHz, based on extensive measurements. 12 TX-RX locations were measured encompassing 5 LOS and 7 NLOS locations spanning 9–38 m. The path loss analysis revealed significant received power even in NLOS scenarios at 6.75 GHz. The PLEs were observed to be small with values below 2 in LOS and close to 2 in NLOS at both 6.75 and 16.95 GHz. The RMS DS was observed to decrease with increasing frequencies in both LOS and NLOS. The RMS AS also captured showed a decreasing trend with frequency and wide angular spreads were observed particularly in NLOS. Frequency-dependent trends for DS, and AS were observed in contrast to some of the constant modeling currently provided in 3GPP specifications.

### v4

@@ III-C RMS Angular Spread
The RMS AS shows the spatial dispersion of arrival or departure angles of MPCs in azimuth and elevation planes. RMS AS is evaluated using a spatial lobe threshold of 10 dB below the strongest pointing direction, using the 3GPP method detailed in [12]. The AS is commonly evaluated using the 3GPP method [12] (circular standard deviation of MPCs) or Fleury’s method (second central moment of the PAS) [29]. A comprehensive comparison between 3GPP and Fleury’s method in computing AS is presented in [29]. Table IV summarizes the AS observed for arrival angles at the RX (ASA) and departure angles at the TX (ASD). In LOS, the ASA is observed as 23.71∘ at 6.75 GHz and 14.32∘ at 16.95 GHz. Wider ASAs of 66.38∘ and 50.40∘ are observed in NLOS at 6.75 GHz and 16.95 GHz, respectively. 3GPP uses a frequency-dependent LOS ASA with overestimated spread, along with a constant NLOS ASA, as outlined in Table 7.5-6 Part-3: Channel Model Parameters for InF and Table 7.5-11: ZSD and ZOD Offset Parameters for InF in [12]. Measured ASAs are observed to decrease with frequency in both LOS and NLOS at 6.75 GHz and 16.95 GHz. Further, we found ASDs to be 35.89∘ in LOS and 45.93∘ in NLOS at 6.75 GHz, while ASDs of 22.91∘ in LOS and 34.77∘ in NLOS were observed at 16.95 GHz. As shown in Fig. 4, the CDFs of empirical data align with these trends, while 3GPP predictions suggest a constant ASD of 39∘ in LOS and 38.9∘ in NLOS from 0.5 GHz to 100 GHz.
Point-data for large-scale spatio-temporal parameters corresponding to each TX-RX location pair are summarized in Table V in concert with Fig. 1. This point-data format enables pooling measurement data from multiple contributors by simply appending rows to the table. Such point-data table can be used to derive traditional statistical models by computing CDFs and scatter plots down each column, such as Fig. 4, while simultaneously preserving the site-specific nature of each measurement location for developing ray-tracing models or applying AI/ML techniques without any additional information other than that denoted in the columns of Table V [26, 29].
@@ IV Conclusion
We presented an in-depth analysis of radio propagation behavior for InF at 6.75 GHz and 16.95 GHz, based on extensive measurements. 12 TX-RX locations were measured, encompassing five LOS and seven NLOS locations spanning 9- 38 m. The path loss analysis revealed significant received power even in NLOS scenarios at 6.75 GHz. The PLEs were observed to be small with values below 2 in LOS and close to 2 in NLOS at both 6.75 and 16.95 GHz. The RMS DS was observed to decrease with increasing frequencies in both LOS and NLOS. The captured RMS AS also showed a decreasing trend with frequency, and wide angular spreads were observed particularly in NLOS. Frequency-dependent trends for DS, and AS were observed in contrast to some of the constant modeling currently provided in 3GPP specifications.